OpenStudy (anonymous):

Let f(x) = 3x - 10. Find the value of f(-3). A. 21 B. -1 C. -19

OpenStudy (anonymous):

C. -19

OpenStudy (anonymous):

c. -19 -3=x so yeah

OpenStudy (anonymous):

Simple! f(3) simply means plug in 3 everywhere you see an x. so f(3) = 3(3)-10. f(3) = 9-10=-19